# On the optimality of dimension truncation error rates for a class of parametric partial differential equations

### v1

@@ 1 Introduction
where $G\!:H_{0}^{1}(D)\to\mathbb{R}$ is a continuous linear quantity of interest and $\mu$ denotes a probability measure over $U$. The measure $\mu$ is typically chosen either as the uniform probability measure over $U=[-1,1]^{\mathbb{N}}$ or as a Gaussian probability measure over $U=\mathbb{R}^{\mathbb{N}}$.
@@
Dimension truncation error rates have so far been primarily analyzed for integration problems. A first upper bound for the dimension truncation rate was established in [11] for a class of elliptic PDEs with an affine parameterization of the diffusion coefficient. This result was subsequently improved in [2] within the broader framework of affine-parametric operator equations. Further studies have addressed dimension truncation for coupled PDE systems arising in optimal control under uncertainty [6], within the periodic setting of uncertainty quantification for numerical integration [9] and kernel interpolation [8], as well as for Bayesian inverse problems governed by PDEs [1, 7]. The analyses in these works rely on Neumann series expansions, a technique that performs well in affine-parametric settings but may yield suboptimal estimates when the parameter dependence is nonlinear. Guth and Kaarnioja [5] subsequently generalized the dimension truncation error analysis to a larger problem class using Taylor series arguments. In particular, the rate obtained for (3) in [5] is $\mathcal{O}(s^{-{2}/{p}+1})$, where $s$ is the truncation dimension and $p$ is related to the decay of the parametric input via
$\bigg\|\prod_{j\geq 1}\frac{\partial^{\nu_{j}}}{\partial y_{j}}a(\cdot,{\boldsymbol{y}})\bigg\|_{L^{\infty}(D)}\leq\Gamma_{|{\boldsymbol{\nu}}|}\prod_{j\geq 1}b_{j}^{\nu_{j}},\quad|{\boldsymbol{\nu}}|:=\sum_{j\geq 1}\nu_{j}<\infty,$
@@ 2 Model problems
$U:=\bigg\{{\boldsymbol{y}}\in\mathbb{R}^{\mathbb{N}}\mid\sum_{j\geq 1}|y_{j}|\|\psi_{j}\|_{L^{\infty}(D)}<\infty\bigg\}.$ (4)
@@
$U:=\bigg\{{\boldsymbol{y}}\in\mathbb{R}^{\mathbb{N}}\mid\sum_{j\geq 1}b_{j}|y_{j}|<\infty\bigg\}.$ (6)
@@ 3 Upper bound
Assume that $(\|\psi_{j}\|_{L^{\infty}(0,1)})\in\ell^{p}(\mathbb{N})$ for some $p\in(0,1)$ such that $\|\psi_{1}\|_{L^{\infty}(0,1)}\geq\|\psi_{2}\|_{L^{\infty}(0,1)}\geq\cdots$. Under the assumptions posed for model problem 1, there holds
$\bigg\|\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}(0,1)}\leq C_{1}s^{-2/p+1},$
where $C_{1}:=\frac{1}{\sqrt{2}}\|f\|_{L^{2}(0,1)}{\rm e}^{\frac{1}{2}\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}(0,1)}^{2}}\big(\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}(0,1)}^{p}\big)^{2/p}$.
@@
$\bigg\|{\frac{\partial}{\partial x}\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|}_{L^{2}(0,1)}^{2}$
@@
$\leq\|f\|_{L^{2}(0,1)}^{2}\int_{0}^{1}\bigg(\int_{U}(\alpha(x,{\boldsymbol{y}})^{-1}-\alpha_{s}(x,{\boldsymbol{y}})^{-1})\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg)^{2}\,{\rm d}x.$
Meanwhile, there holds
$\bigg\|\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{L^{2}(0,1)}^{2}$
@@
$\leq\|f\|_{L^{2}(0,1)}^{2}\int_{0}^{1}\bigg(\int_{U}(\alpha(\xi,{\boldsymbol{y}})^{-1}-\alpha_{s}(\xi,{\boldsymbol{y}})^{-1})\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg)^{2}\,{\rm d}\xi.$
This means that the $H^{1}$-error satisfies
$\bigg\|\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}(0,1)}$
$\leq\sqrt{2}\|f\|_{L^{2}(0,1)}\sqrt{\int_{0}^{1}\bigg(\int_{U}(\alpha(x,{\boldsymbol{y}})^{-1}-\alpha_{s}(x,{\boldsymbol{y}})^{-1})\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg)^{2}\,{\rm d}x}.$
@@
where we used the identity $\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}{\rm e}^{by}{\rm e}^{-\frac{1}{2}y^{2}}\,{\rm d}y={\rm e}^{\frac{1}{2}b^{2}}$ for $b\in\mathbb{R}$. Substituting this into the integral and using ${\rm e}^{t}-1\leq t{\rm e}^{t}$ for $t\geq 0$ together with the $\ell^{p}$-summability of $(\lVert\psi_{j}\rVert_{L^{\infty}})$ gives
$\bigg|\int_{U}(\alpha(x,{\boldsymbol{y}})^{-1}-\alpha_{s}(x,{\boldsymbol{y}})^{-1})\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg|\leq\frac{1}{2}{\rm e}^{\frac{1}{2}\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}(0,1)}^{2}}\sum_{j>s}\|\psi_{j}\|_{L^{\infty}(0,1)}^{2}.$
As a consequence of Stechkin’s lemma (cf., e.g., [10, Lemma 3.3]), there holds
$\sum_{j>s}\|\psi_{j}\|_{L^{\infty}(0,1)}^{2}\leq s^{-2/p+1}\bigg(\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}(0,1)}^{p}\bigg)^{2/p},$
@@
$\bigg\|{\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})}\bigg\|_{H^{1}(0,1)}$
$\leq\frac{1}{\sqrt{2}}\|f\|_{L^{2}(0,1)}{\rm e}^{\frac{1}{2}\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}(0,1)}^{2}}\bigg(\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}(0,1)}^{p}\bigg)^{2/p}s^{-2/p+1},$
@@
where $C_{2}:=\|\widetilde{w}\|_{H^{1}(D)}\frac{1}{2}{\rm e}^{\frac{1}{2}\sum_{j=1}^{\infty}b_{j}^{2}}\bigg(\sum_{j=1}^{\infty}b_{j}^{p}\bigg)^{2/p}$.
@@ 4 Lower bound
$\bigg\|\int_{U}(u(\cdot,{\boldsymbol{y}})-u_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}(0,1)}\geq\frac{C_{f}}{2\theta-1}s^{-2\theta+1},$
@@
$\int_{U}(u(x,{\boldsymbol{y}})-u_{s}(x,{\boldsymbol{y}}))\,\mu_{G}(\mathrm{d}{\boldsymbol{y}})$
@@
$\bigg\|\int_{U}(u(\cdot,{\boldsymbol{y}})-u_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}}^{2}$
@@
$\bigg\|\int_{U}(u(\cdot,{\boldsymbol{y}})-u_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}(0,1)}^{2}$ $\geq\int_{0}^{1}\bigg(\int_{x}^{1}f(z)\,{\rm d}z\bigg)^{2}\frac{1}{4}\Big(\sum_{j>s}\psi_{j}(x)^{2}\Big)^{2}\,\mathrm{d}x$
@@
$\bigg\|{\int_{U}(u(\cdot,{\boldsymbol{y}})-u_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})}\bigg\|_{H^{1}(0,1)}\geq\frac{C_{f}}{2\theta-1}\,s^{-2\theta+1}.$
@@
As $\theta\searrow 1/p$, the lower bound exhibits the same algebraic rate as the upper bound, up to multiplicative constants, confirming the asymptotic sharpness of the estimate.
Lemma 4 (Dimension truncation upper bound for model problem 2).

### v2

@@ 1 Introduction
where $G\!:H_{0}^{1}(D)\to\mathbb{R}$ is a continuous linear quantity of interest and $\mu$ denotes a probability measure over $U$. By $H_{0}^{1}(D)$ we denote the subspace of $H^{1}(D)$ with zero trace on $\partial D$. The measure $\mu$ is typically chosen either as the uniform probability measure over $U=[-1,1]^{\mathbb{N}}$ or as a Gaussian probability measure over $U=\mathbb{R}^{\mathbb{N}}$.
@@
Dimension truncation error rates have so far been primarily analyzed for integration problems. A first upper bound for the dimension truncation rate was established in [12] for a class of elliptic PDEs with an affine parameterization of the diffusion coefficient. This result was subsequently improved in [2] within the broader framework of affine-parametric operator equations. Further studies have addressed dimension truncation for coupled PDE systems arising in optimal control under uncertainty [6], within the periodic setting of uncertainty quantification for numerical integration [9] and kernel interpolation [8], as well as for Bayesian inverse problems governed by PDEs [1, 7]. The analyses in these works rely on Neumann series expansions (see, e.g., [10, Example 4.5]), a technique that performs well in affine-parametric settings but may yield suboptimal estimates when the parameter dependence is nonlinear. Guth and Kaarnioja [5] subsequently generalized the dimension truncation error analysis to a larger problem class using Taylor series arguments. In particular, the rate obtained for (3) in [5] is $\mathcal{O}(s^{-{2}/{p}+1})$, where $s$ is the truncation dimension and $p$ is related to the decay of the parametric input via
$\bigg\|\prod_{j\geq 1}\frac{\partial^{\nu_{j}}}{\partial y_{j}^{\nu_{j}}}a(\cdot,{\boldsymbol{y}})\bigg\|_{L^{\infty}(D)}\leq\Gamma_{|{\boldsymbol{\nu}}|}\prod_{j\geq 1}b_{j}^{\nu_{j}},\quad|{\boldsymbol{\nu}}|:=\sum_{j\geq 1}\nu_{j}<\infty,$
@@ 2 Model problems
$U:=\bigg\{{\boldsymbol{y}}\in\mathbb{R}^{\mathbb{N}}:\sum_{j\geq 1}|y_{j}|\|\psi_{j}\|_{L^{\infty}(D)}<\infty\bigg\}.$ (4)
@@
$U:=\bigg\{{\boldsymbol{y}}\in\mathbb{R}^{\mathbb{N}}:\sum_{j\geq 1}b_{j}|y_{j}|<\infty\bigg\}.$ (6)
@@ 3 Upper bound
Assume that $(\|\psi_{j}\|_{L^{\infty}((0,1))})_{j\geq 1}\in\ell^{p}(\mathbb{N})$ for some $p\in(0,1)$ such that $\|\psi_{1}\|_{L^{\infty}((0,1))}\allowbreak\geq\|\psi_{2}\|_{L^{\infty}((0,1))}\geq\cdots$. Under the assumptions posed for model problem 1, there holds
$\bigg\|\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}((0,1))}\leq C_{1}s^{-2/p+1},$
where $C_{1}:=\frac{1}{\sqrt{2}}\|f\|_{L^{2}((0,1))}{\rm e}^{\frac{1}{2}\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}((0,1))}^{2}}\big(\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}((0,1))}^{p}\big)^{2/p}$.
@@
$\bigg\|{\frac{\partial}{\partial x}\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|}_{L^{2}((0,1))}^{2}$
@@
$\leq\|f\|_{L^{2}((0,1))}^{2}\int_{0}^{1}\bigg(\int_{U}(\alpha(x,{\boldsymbol{y}})^{-1}-\alpha_{s}(x,{\boldsymbol{y}})^{-1})\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg)^{2}\,{\rm d}x,$
where we interchange the order of differentiation and integration in the first step. Then, after an application of the Cauchy–Schwarz inequality, the first factor can be bounded uniformly in $x$ by enlarging the domain of integration. Meanwhile, there holds
$\bigg\|\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{L^{2}((0,1))}^{2}$
@@
$\leq\|f\|_{L^{2}((0,1))}^{2}\int_{0}^{1}\bigg(\int_{U}(\alpha(\xi,{\boldsymbol{y}})^{-1}-\alpha_{s}(\xi,{\boldsymbol{y}})^{-1})\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg)^{2}\,{\rm d}\xi,$
where we applied the Cauchy–Schwarz inequality and bounded the expression from above by enlarging the domains of integration. This means that the $H^{1}$-error satisfies
$\bigg\|\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}((0,1))}$
$\leq\sqrt{2}\|f\|_{L^{2}((0,1))}\sqrt{\int_{0}^{1}\bigg(\int_{U}(\alpha(x,{\boldsymbol{y}})^{-1}-\alpha_{s}(x,{\boldsymbol{y}})^{-1})\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg)^{2}\,{\rm d}x}.$
@@
where we used the identity $\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}{\rm e}^{by}{\rm e}^{-\frac{1}{2}y^{2}}\,{\rm d}y={\rm e}^{\frac{1}{2}b^{2}}$ for $b\in\mathbb{R}$. Then, using ${\rm e}^{t}-1\leq t{\rm e}^{t}$ for $t\geq 0$ together with the $\ell^{p}$-summability of $(\lVert\psi_{j}\rVert_{L^{\infty}((0,1))})_{j\geq 1}$ gives
$\bigg|\int_{U}(\alpha(x,{\boldsymbol{y}})^{-1}-\alpha_{s}(x,{\boldsymbol{y}})^{-1})\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg|\leq\frac{1}{2}{\rm e}^{\frac{1}{2}\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}((0,1))}^{2}}\sum_{j>s}\|\psi_{j}\|_{L^{\infty}((0,1))}^{2}.$
As a consequence of Stechkin’s lemma (cf., e.g., [11, Lemma 3.3]), there holds
$\sum_{j>s}\|\psi_{j}\|_{L^{\infty}((0,1))}^{2}\leq s^{-2/p+1}\bigg(\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}((0,1))}^{p}\bigg)^{2/p},$
@@
$\bigg\|{\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})}\bigg\|_{H^{1}((0,1))}$
$\leq\frac{1}{\sqrt{2}}\|f\|_{L^{2}((0,1))}{\rm e}^{\frac{1}{2}\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}((0,1))}^{2}}\bigg(\sum_{j=1}^{\infty}\|\psi_{j}\|_{L^{\infty}((0,1))}^{p}\bigg)^{2/p}s^{-2/p+1},$
@@
where $C_{2}:=\|\widetilde{w}\|_{H^{1}(D)}\frac{1}{2}{\rm e}^{\frac{1}{2}\sum_{j=1}^{\infty}b_{j}^{2}}\big(\sum_{j=1}^{\infty}b_{j}^{p}\big)^{2/p}$.
@@ 4 Lower bound
$\bigg\|\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}((0,1))}\geq\frac{C_{f}}{2\theta-1}s^{-2\theta+1},$
@@
$\int_{U}(v(x,{\boldsymbol{y}})-v_{s}(x,{\boldsymbol{y}}))\,\mu_{G}(\mathrm{d}{\boldsymbol{y}})$
@@
$\bigg\|\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}}^{2}$
@@
$\bigg\|\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}((0,1))}^{2}$ $\geq\int_{0}^{1}\bigg(\int_{x}^{1}f(z)\,{\rm d}z\bigg)^{2}\frac{1}{4}\Big(\sum_{j>s}\psi_{j}(x)^{2}\Big)^{2}\,\mathrm{d}x$
@@
$\bigg\|{\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})}\bigg\|_{H^{1}((0,1))}\geq\frac{C_{f}}{2\theta-1}\,s^{-2\theta+1}.$
@@
As $\theta\searrow 1/p$, the lower bound exhibits the same algebraic rate as the upper bound, up to multiplicative constants, confirming the asymptotic sharpness of the estimate. Consider for example a sequence $(\psi_{j})_{j\geq 1}$ satisfying $\|\psi_{j}\|_{L^{\infty}((0,1))}=cj^{-\theta}$ with $\theta>1$. By Lemma 3, there holds
$\bigg\|\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}((0,1))}\geq\frac{C_{f}}{2\theta-1}s^{-2\theta+1}.$
Meanwhile, $(\|\psi_{j}\|_{L^{\infty}((0,1))})_{j\geq 1}\in\ell^{p}(\mathbb{N})$ for all $p\in(\frac{1}{\theta},1)$. By Lemma 1 and choosing $p=\frac{2}{2\theta-\varepsilon}$ for arbitrary $\varepsilon>0$, there holds
$\bigg\|\int_{U}(v(\cdot,{\boldsymbol{y}})-v_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}((0,1))}\leq C_{1}s^{-\frac{2}{p}+1}=C_{1}s^{-2\theta+1+\varepsilon}.$
Lemma 4 (Dimension truncation lower bound for model problem 2).
@@
The asymptotic sharpness follows by the same reasoning as above: considering $b_{j}=cj^{-\theta}$ with $\theta>1$, Lemmas 2 and 4 imply that
$\bigg\|\int_{U}(w(\cdot,{\boldsymbol{y}})-w_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}(D)}\geq\frac{C_{\widetilde{w}}}{2\theta-1}s^{-2\theta+1}$
and
$\bigg\|\int_{U}(w(\cdot,{\boldsymbol{y}})-w_{s}(\cdot,{\boldsymbol{y}}))\,\mu_{G}({\rm d}{\boldsymbol{y}})\bigg\|_{H^{1}(D)}\leq C_{2}s^{-2\theta+1+\varepsilon}$
for arbitrary $\varepsilon>0$.